What volume of 0.750M KCl solution is required to react exactly with 1325.0mL of 1.352 M Pb(NO3)2 solution
1
Expert's answer
2020-03-18T12:55:54-0400

"2KCl +Pb(NO_3)_2\rightarrow PbCl_2+ 2KNO_3"

"1325 mL \times \frac{1L}{1000mL}= 1.325L"

moles of "(Pb(NO_3)_2)= 1.325 L \times \frac{1.352 mol}{1L}=1.791 moles"

"1.791 moles Pb(NO_3)_2\times \frac{2 mol KCl}{1 mole Pb(NO_3)_2}=3.583 mol KCl"

find volume of KCl solution:

"V=\frac{n}{c} = \frac{3.583 mol KCl}{0.750 \frac{mol}{L}}=4.777 L=4777 mL"